OpenStudy (liddoislander):

What is the percent of change from $315.00 to $220.00?

OpenStudy (amoodarya):

315-220 --------- * 100 315

OpenStudy (liddoislander):

these are the choices ; A.30% decrease B.30% increase C.43% decrease D.43% increase

OpenStudy (liddoislander):

@amoodarya

OpenStudy (amoodarya):

FIRST decreases because 315 - ( )=220 then simplify the ratio

OpenStudy (amoodarya):

315-220 95 19 --------- * 100 =-------*100 =----- *100 =0.301*100=30.1 % 315 315 63

OpenStudy (liddoislander):

so A. 30% decreases ?